# Efficient Generative Model Training via Embedded Representation Warmup

### v2

@@ Abstract
Generative models face a fundamental challenge: they must simultaneously learn high-level semantic concepts (what to generate) and low-level synthesis details (how to generate it). Conventional end-to-end training entangles these distinct, and often conflicting objectives, leading to a complex and inefficient optimization process. We argue that explicitly decoupling these tasks is key to unlocking more effective and efficient generative modeling. To this end, we propose Embedded Representation Warmup (ERW), a principled two-phase training framework. The first phase is dedicated to building a robust semantic foundation by aligning the early layers of a diffusion model with a powerful pretrained encoder. This provides a strong representational prior, allowing the second phase—generative full training with alignment loss to refine the representation—to focus its resources on high-fidelity synthesis. By separating the learning of "what" from "how," ERW fundamentally simplifies the optimization landscape. Our analysis confirms that this efficacy stems from functionally specializing the model’s early layers for representation. Empirically, our framework achieves up to a 11.5$\times$ speedup in training iterations to reach FID=1.41 compared to single-phase methods like REPA (Yu et al., 2024). Code is available at https://github.com/LINs-lab/ERW.
@@ 1 Introduction
Validation. Extensive experiments demonstrate that our decoupling strategy yields substantial benefits. ERW achieves up to an $\mathbf{11.5\times}$ speedup in training iterations compared to single-phase methods like REPA while achieving FID=1.41. The warmup phase requires only a fraction of the total training cost, making our approach highly practical for real-world applications.
@@
    We formalize the optimization entanglement in generative models as a conflict between semantic understanding and visual synthesis, and propose a conceptual decomposition of the diffusion model into functionally specialized L2R and R2G circuits.
@@ 2 Related Work
Numerous studies confirm that diffusion models implicitly learn powerful, classifier-like semantic features (Yang & Wang, 2023; Li et al., 2023a; Xiang et al., 2023), a phenomenon some works have deconstructed this phenomenon for self-supervised learning (Chen et al., 2024). An alternative strategy enhances internal representation learning by fusing diffusion objectives with auxiliary self-supervision losses, exemplified by MAGE (Li et al., 2023c) and MaskDiT (Zheng et al., 2024), which draw inspiration from contrastive learning (Chen et al., 2020a; He et al., 2020) and masked autoencoding (He et al., 2022). However, these approaches require careful balancing of competing objectives. ERW sidesteps these complexities by directly injecting mature semantic priors via dedicated warmup, freeing the model to focus purely on high-fidelity synthesis while achieving training efficiency comparable to contemporary acceleration methods.
3 Hypothesis of Latent Diffusion Models
In this section, we propose a three-stage diffusion circuit—Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G)—as a structured framework for more effective and interpretable generative modeling. Our guiding hypothesis is that conventional diffusion methods entangle representation learning and generative decoding within the denoising process, burdening the network with two inherently distinct tasks. By explicitly isolating the semantic processing (L2R) from the generative inversion (R2G), our approach promotes modularity and simplifies the overall optimization landscape. This theoretical perspective also clarifies how different regions of the diffusion backbone specialize in semantic vs. generative aspects, offering a fresh viewpoint on the internal mechanics of diffusion models.
@@ 3.1 Preliminaries
Consider a data sample $\mathbf{x}\sim p(\mathbf{x})$ (e.g., an image) and let the encoder $\mathcal{H}_{\bm{\theta}}(\mathbf{x})$ map it to its latent representation denoted as $\mathbf{z}\in\mathcal{Z}$. Given standard Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, SiT defines a forward process in the latent space, parameterized by continuous time $t\in[0,1]$:
$\mathbf{z}_{t}=\alpha_{t}\,\mathbf{z}\,+\,\sigma_{t}\,\bm{\epsilon}\,,$ (1)
@@
This construction implies that $\mathbf{z}_{0}=\mathbf{z}$ (the clean latent representation) and $\mathbf{z}_{1}=\bm{\epsilon}$ (pure noise). Under mild conditions (Albergo et al., 2023), the sequence $\{\mathbf{z}_{t}\}$ forms a stochastic interpolant that smoothly transitions between data and noise in the latent space.
@@
$\dot{\mathbf{z}}_{t}=\dot{\alpha}_{t}\,\mathbf{z}\,+\,\dot{\sigma}_{t}\,\bm{\epsilon}\,.$ (3)
@@
$\dot{\mathbf{z}}_{t}=\mathbf{v}(\mathbf{z}_{t},t)\,,$ (4)
where $\mathbf{v}(\mathbf{z}_{t},t)$ is defined as the conditional expectation of $\dot{\mathbf{z}}_{t}$ given $\mathbf{z}_{t}$. A neural network $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ is then trained to approximate $\mathbf{v}(\mathbf{z},t)$ by minimizing:
$\mathcal{L}_{\mathrm{velocity}}(\bm{\theta})\;=\;{\mathbb{E}}_{\mathbf{z},\bm{\epsilon},t}\left[\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t)-\Bigl{(}\dot{\alpha}_{t}\,\mathbf{z}+\dot{\sigma}_{t}\,\bm{\epsilon}\Bigr{)}\right\rVert^{2}\right]\,.$ (5)
Successfully learning $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ enables one to integrate the reverse-time ordinary differential equation (ODE) (Song et al., 2020), thereby mapping noise samples back to coherent latent representations.
3.2 Three-Stage Diffusion Circuit
Figure 2: Illustration of the Circuits in Hypothesis. From left to right, we first apply a VAE encoder (the Pixel-to-Latent or (P2L) stage) to map high-dimensional inputs to a compressed latent space. We then perform latent diffusion on these codes, dividing the backbone into two subregions: the Latent-to-Representation (L2R) region that capture and refine semantic features, and the Representation-to-Generation (R2G) region that decode the learned representation into final outputs.
@@ 3.2 Three-Stage Diffusion Circuit
Here, the Variational Autoencoder (VAE) encoder $\mathcal{H}_{\bm{\theta}}$ compresses high-dimensional pixel data $\mathbf{x}$ into a lower-dimensional latent $\mathbf{z}$. The diffusion model components $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ and $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ share parameters in diffusion model backbone but serve distinct functions: $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ extracts semantic representations from noisy latent codes (L2R), while $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ transforms these representations into refined latent outputs (R2G). During inference, a VAE decoder $\mathcal{D}_{\bm{\theta}}$ would convert these refined latents back to pixel space. This decomposition into distinct processing stages provides a more transparent view of how latent diffusion models function and support the idea from latent space compression in Saharia et al. (2022) and the diffusion model training via representation alignment in REPA Yu et al. (2024).
@@
The diffusion model’s objective is to learn the score function $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})$, which represents the gradient of the log probability density of the data distribution. Our conceptual motivation is that this score function can be orthogonally decomposed into representation and generation components under certain idealizing assumptions (see details in Appendix A):
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})}_{\mathrm{Representation}}+\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})}_{\mathrm{Generation}}\,.$ (6)
This decomposition permits us to partition the overall diffusion objective into two nearly orthogonal terms, focusing respectively on (i) extracting meaningful features from corrupted inputs and (ii) translating those features into coherent outputs. While this decomposition is an idealization, as the components are not perfectly orthogonal in a real network and the intermediate representation $\mathbf{r}_{t}$ is not a perfect sufficient statistic, it provides a useful lens for our method. Such a structure naturally aligns with the three-stage diffusion circuit we will soon discuss, clarifying the distinct roles played by different network regions.
@@
Before performing the denoising process in the high-dimensional pixel domain—where noise may obscure semantic cues—many contemporary methods Saharia et al. (2022); Ho et al. (2020); Dhariwal & Nichol (2021) compress images into a more tractable latent space:
@@
Figure 3: Selected Samples on ImageNet $\times 256$ 256. Images generated by the SiT-XL/2 + REPA + ERW model using Classifier-Free Guidance (CFG) with a scale of $w=1.62$ under 350 epochs.
@@
Given a noisy latent $\mathbf{z}_{t}$ from Eq. (7), the model initially extracts a semantic representation $\mathbf{r}_{t}$ using the mapping $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$.
@@
This step corresponds to estimating $\nabla\log p(\mathbf{r}_{t})$ within our conceptual velocity loss decomposition (where we separate representation learning from generation, as detailed in  Appendix A), sometimes termed the Representation Score Song et al. (2020). Intuitively, the model should discern salient patterns (e.g., object shapes, style characteristics, or conditioning signals) before denoising. Under the sufficient statistic assumption (see Assump. 1), the representation $\mathbf{r}_{t}$ effectively captures the essential information from the latent $\mathbf{z}_{t}$, allowing us to formulate the representation learning process as an independent objective:
$\mathcal{L}_{\mathrm{L2R}}\;=\;\mathbb{E}_{t,\mathbf{z}_{t}}\bigl{[}\|\,\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\,\|^{2}\bigr{]}\,,$ (9)
where $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ represents the parameterized representation extraction module of the network. By explicitly decoupling the objective for semantic feature extraction from that of generative refinement, the model is guided to learn robust high-level representations and ensures that the early layers focus on capturing essential semantic features.
@@
In the final part of each reverse diffusion update in (3), known as the R2G stage, the extracted semantic representation is transformed into an updated latent with reduced noise:
@@
This output serves the same purpose as the $\mathbf{z}^{\prime}$ term introduced in our framework’s third Eq. (3), but is specifically defined for the discrete time step $t-\Delta t$ in the continuous-time diffusion process. In the decomposed view, this step aligns with the Generation Score term, $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$, detailed in  Appendix A:
$\mathcal{L}_{\mathrm{R2G}}\!=\!\mathbb{E}_{t,\mathbf{z}_{t}}\Big{[}\left\lVert\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\!\mid\!\mathbf{r}_{t})\right\rVert^{2}\Big{]}\,,$ (11)
where $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ represents the generation module’s score estimator, transforming representations back into the latent space for denoising. Injecting the semantic representation $\mathbf{r}_{t}$ into a cleaner latent $\mathbf{z}_{t-\Delta t}$—which is significantly less noisy than $\mathbf{z}_{t}$—ensures that abstract semantic features are effectively transformed into the precise latent elements required for content generation. Meanwhile, the cross term between L2R and R2G (also elaborated in  Appendix A) remains small under the assumption that these two gradients are sufficiently orthogonal, mitigating destructive interference.
@@
Hence, every time step naturally splits into (i) L2R for refining the representation and (ii) R2G for synthesizing an updated latent. This loop neatly implements the principle of “first representation, then generation”. Empirically, prior work (Yu et al., 2024; Xiang et al., 2023) confirms that early layers of the diffusion model predominantly focus on representation extraction, whereas later layers emphasize generative refinement. Consequently, the staged design mirrors the reverse-time diffusion trajectory, concluding in a final latent $\mathbf{z}_{0}$ that is decoded via $\mathcal{D}_{\bm{\theta}}$ to yield the synthesized output $\mathbf{x}_{0}$.
4 Embedded Representation Warmup: A Two-Phase Training Framework
Diffusion models generate high-fidelity images (Ho et al., 2020; Dhariwal & Nichol, 2021), yet their training is impeded by the simultaneous challenges of learning robust semantic features and performing generative denoising (Song et al., 2020). This coupling creates an inherently complex optimization landscape that demands prolonged training periods.
@@ 4 Embedded Representation Warmup: A Two-Phase Training Framework
In this section, we present Embedded Representation Warmup (ERW), a framework that strategically decouples these challenges into two distinct phases. In Phase 1, we initialize the early layers of the diffusion model with high-quality semantic features from pretrained models; in Phase 2, we transition to main diffusion training with a gradually diminishing representation alignment term, allowing the model to increasingly focus on generation. This separation addresses the inefficiency of joint optimization in conventional diffusion models and significantly enhances training efficiency.
4.1 Phase 1: Representation Warmup Stage
@@
To alleviate the burden of learning semantic features from scratch, we begin with a dedicated warmup stage. During this phase, the model’s L2R circuit is initialized to align with semantically rich features extracted from a pretrained representation model (e.g., DINOv2, MAE, or CLIP).
Let $\mathcal{H}_{\mathbf{\bm{\theta}}}(\mathbf{x})$ denote an encoder that maps an image $\mathbf{x}\in\mathcal{X}$ to its latent representation $\mathbf{z}\in\mathcal{Z}$, and let $\mathbf{f}_{\mathrm{rep}}:\mathcal{X}\to\mathcal{R}$ be a high-quality pretrained representation model. Our objective is to learn a mapping $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}$ such that its output closely aligns with $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$.
Specifically, we define the warmup loss as:
$\mathcal{L}_{\mathrm{warmup}}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\Bigl{[}\bigl{\|}\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\bm{\theta}}}(\mathbf{x}))-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\bigr{\|}_{2}^{2}\Bigr{]}$ (13)
By minimizing Eq. (13), we ensure that the early layers of the diffusion model, denoted $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}$, produce features on clean latents that are well-aligned with the pretrained representation model. This alignment mitigates the need for the diffusion model to learn complex semantic features purely from noisy inputs; instead, subsequent layers can operate on features that are already semantically meaningful.
4.2 Phase 2: Generative Training with Decaying Representation Guidance
@@
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{diffusion}}+\lambda(t)\cdot\mathcal{L}_{\mathrm{align}}$ (14)
Here, $\mathcal{L}_{\mathrm{diffusion}}$ denotes the conventional diffusion loss (e.g., the velocity prediction loss as defined in Eq. (5)), $\mathcal{L}_{\mathrm{align}}$ is the representation alignment loss, and $\lambda(t)$ is a time-dependent weight that modulates the impact of alignment during training.
In this phase, the alignment loss $\mathcal{L}_{\mathrm{align}}$ encourages the diffusion model’s hidden representations for noisy inputs to align with the external encoder’s features for clean inputs, following REPA. A trainable projection head $\mathcal{T}_{\mathbf{\bm{\theta}}}$ maps the intermediate features to the representation space:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{t,\mathbf{x},\mathbf{\epsilon}}\Biggl{[}\Bigl{\|}\mathcal{T}_{\mathbf{\bm{\theta}}}\Bigl{(}\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\Bigr{)}-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\Bigr{\|}_{2}^{2}\Biggr{]}$ (15)
where $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)$ is the intermediate representation from the early layers of the model given noisy latent $\mathbf{z}_{t}$.
The time-dependent alignment weight follows an exponential decay schedule:
$\lambda(t)=c_{0}\exp\!\left(-\frac{t}{\tau}\right)$ (16)
where $c_{0}$ is the initial weight of the alignment loss and $\tau$ is a hyperparameter controlling the decay rate. This schedule provides strong semantic guidance during early training—when representation learning is most critical—and gradually reduces its influence, allowing the network to increasingly focus on optimizing generative capabilities.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Validation. Extensive experiments demonstrate that our decoupling strategy yields substantial benefits. ERW achieves up to an $\mathbf{11.5\times}$ training speedup to reach a comparable FID score in 350 epochs compared to single-phase methods like REPA while achieving $\mathbf{FID=1.41}$. The warmup phase requires only a fraction of the total training cost, making our approach highly practical for real-world applications.
@@
    We formalize the optimization entanglement in generative models as of semantic understanding and visual synthesis, and propose a conceptual decomposition of the diffusion model into functionally specialized L2R and R2G circuits.
@@ 2 Related Work
Numerous studies confirm that diffusion models implicitly learn powerful, classifier-like semantic features (Yang & Wang, 2023; Li et al., 2023a; Xiang et al., 2023), a phenomenon some works have deconstructed this phenomenon for self-supervised learning (Chen et al., 2024). An alternative strategy enhances internal representation learning by fusing diffusion objectives with auxiliary self-supervision losses, exemplified by MAGE (Li et al., 2023c) and MaskDiT (Zheng et al., 2024), which draw inspiration from contrastive learning (Chen et al., 2020a; He et al., 2020) and masked autoencoding (He et al., 2022). However, these approaches require careful balancing of competing objectives. ERW sidesteps these complexities by directly injecting mature semantic priors via dedicated warmup, freeing the model to focus purely on high-fidelity synthesis while achieving efficiency comparable to contemporary methods.
3 From Functional Specialization to Decoupled Training in Latent Diffusion
In this section, we adopt a three-stage view of latent diffusion—Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G)— as a functional perspective that facilitates decoupled training. P2L provides compressed latents as a precondition, while L2R and R2G capture the predominant (but not exclusive) roles of early and late layers in semantic processing and generative refinement. The separation is heuristic and approximate—roles overlap and are not strictly orthogonal—but it is sufficient to decouple training objectives in practice. This view underpins our two-phase framework.
@@ 3.1 Preliminaries
Consider a data sample $\mathbf{x}\sim p(\mathbf{x})$ (e.g., an image) and let the encoder $\mathcal{H}_{\bm{\theta}}(\mathbf{x})$ map it to its latent representation denoted as $\mathbf{z}_{0}\in\mathcal{Z}$. Given standard Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, SiT defines a forward process in the latent space, parameterized by continuous time $t\in[0,1]$:
$\mathbf{z}_{t}=\alpha_{t}\,\mathbf{z}_{0}\,+\,\sigma_{t}\,\bm{\epsilon}\,$, (1)
@@
This construction implies that at $t=0$ we recover the clean latent $\mathbf{z}_{0}$, and at $t=1$ we have pure noise $\mathbf{z}_{1}=\bm{\epsilon}$. Under mild conditions (Albergo et al., 2023), the sequence $\{\mathbf{z}_{t}\}$ forms a stochastic interpolant that smoothly transitions between data and noise in the latent space.
@@
$\dot{\mathbf{z}}_{t}=\dot{\alpha}_{t}\,\mathbf{z}_{0}\,+\,\dot{\sigma}_{t}\,\bm{\epsilon}\,.$ (3)
@@
$\dot{\mathbf{z}}_{t}=\mathbf{F}(\mathbf{z}_{t},t)\,$, (4)
where $\mathbf{F}(\mathbf{z}_{t},t)$ is defined as the conditional expectation of $\dot{\mathbf{z}}_{t}$ given $\mathbf{z}_{t}$. A neural network $\mathbf{F}_{\bm{\theta}}(\mathbf{z},t)$ is then trained to approximate $\mathbf{F}(\mathbf{z},t)$ by minimizing:
$\mathcal{L}_{\mathrm{diffusion}}(\bm{\theta})\;=\;{\mathbb{E}}_{\mathbf{z}_{0},\bm{\epsilon},t}\left[\left\lVert\mathbf{F}_{\bm{\theta}}(\mathbf{z}_{t},t)-\Bigl(\dot{\alpha}_{t}\,\mathbf{z}_{0}+\dot{\sigma}_{t}\,\bm{\epsilon}\Bigr)\right\rVert^{2}\right]\,.$ (5)
Learning $\mathbf{F}_{\bm{\theta}}(\mathbf{z},t)$ enables one to integrate the reverse-time ordinary differential equation (ODE) (Song et al., 2020), thereby mapping noise samples back to coherent latent representations.
3.2 A Functional Circuit Perspective for Decoupled Training
Figure 2: Functional circuit for decoupled training. P2L provides a compression precondition; early layers predominantly serve Latent-to-Representation (L2R, semantic inference), and later layers predominantly serve Representation-to-Generation (R2G, synthesis). Roles overlap in practice; we employ this perspective to organize objectives and reduce optimization entanglement.
@@ 3.2 A Functional Circuit Perspective for Decoupled Training
Here, $\mathcal{H}_{\bm{\theta}}$ is a VAE encoder that compresses pixels to latents; $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ and $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ are two overlapping functional roles implemented within the shared diffusion backbone. A VAE decoder $\mathcal{D}_{\bm{\theta}}$ maps refined latents back to pixels at the end.
@@
Grounded in the augmented probability view, Appendix B (Thm. 1) gives an exact decomposition of the joint conditional score:
[…]
This provides a principled rationale for separating optimization into representation inference (L2R) and conditional generation (R2G). In practice, we shape these two components using surrogate losses: the standard diffusion objective in Eq. (5) for generation and the alignment objective in Eq. (14) for representation; see Appendix B for details.
@@
Before performing the denoising process in the high-dimensional pixel domain—where noise may obscure semantic cues—many methods Saharia et al. (2022); Ho et al. (2020); Dhariwal & Nichol (2021) compress images into a more tractable latent space:
@@
Figure 3: Selected Samples on ImageNet $256\times 256$. Images generated by the SiT-XL/2 + REPA + ERW model using Classifier-Free Guidance (CFG) with a scale of $w=1.62$ under 350 epochs.
@@
Given a noisy latent $\mathbf{z}_{t}$ from the forward process (Eq. (1)), the model initially extracts a semantic representation $\mathbf{r}_{t}$ using the mapping $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$.
@@
This step corresponds to the Representation Inference Score, i.e., estimating $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ in the augmented conditional view (Thm. 1; see also  Appendix B). Intuitively, the model should discern salient patterns (e.g., object shapes, style characteristics, or conditioning signals) before denoising. Under the sufficient statistic assumption (see Assump. 1), the representation $\mathbf{r}_{t}$ effectively captures the essential information from the latent $\mathbf{z}_{t}$. The true representation score available one could consider the idealized regression objective
$\min_{\mathcal{R}}\;\mathbb{E}_{t,\mathbf{z}_{t}}\bigl[\|\,\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)\,\|^{2}\bigr]\,.$ (9)
In practice, we do not access this score; instead we employ surrogate alignment losses: the clean-latent warmup in Eq. (13) and the noisy-input alignment term in Eq. (14). By explicitly decoupling the objective for semantic feature extraction from that of generative refinement, the model is guided to learn representations and ensures that the early layers focus on capturing semantic features.
@@
In the final phase of each reverse diffusion update in (3), known as the R2G stage, the extracted semantic representation is transformed into an updated latent with reduced noise:
@@
This output serves the same purpose as the $\mathbf{z}^{\prime}$ term introduced, but is specifically defined for the discrete time step $t-\Delta t$ in the continuous-time diffusion process. In the decomposition, this step aligns with the Conditional Generation Score component. For the rigorous joint-conditional view, see Thm. 1. The conditional generation score available one could consider the idealized regression objective
$\min_{\mathcal{G}}\;\mathbb{E}_{t,\mathbf{z}_{t}}\Big[\left\lVert\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\!\mid\!\mathbf{z}_{t},\mathbf{r}_{t},t)\right\rVert^{2}\Big]\,.$ (11)
In practice, we instead rely on the standard diffusion objective in Eq. (5) (and its Phase 2 combination in Eq. (14)) to shape the generation component while using the learned representations as guidance. Injecting the semantic representation $\mathbf{r}_{t}$ into a cleaner latent $\mathbf{z}_{t-\Delta t}$—which is significantly less noisy than $\mathbf{z}_{t}$—ensures that abstract semantic features are effectively transformed into the precise latent elements required for content generation. Meanwhile, the cross interaction between L2R and R2G (also discussed in  Appendix B) is empirically small when the two gradients are sufficiently separated in function, helping to mitigate destructive interference.
@@
Hence, every time step naturally splits into (i) L2R for refining the representation and (ii) R2G for synthesizing an updated latent. This loop neatly implements the principle of "first representation, then generation". Empirically, prior work (Yu et al., 2024; Xiang et al., 2023) confirms that early layers of the diffusion model predominantly focus on representation extraction, whereas later layers emphasize generative refinement. Consequently, the staged design mirrors the reverse-time diffusion trajectory, concluding in a final latent $\mathbf{z}_{0}$ that is decoded via $\mathcal{D}_{\bm{\theta}}$ to yield the synthesized output $\mathbf{x}_{0}$.
3.3 Embedded Representation Warmup: Training with Two Phases
@@
Guided by the circuit view and our augmented-space analysis, we present Embedded Representation Warmup (ERW), a framework that strategically decouples training into two phases. In Phase 1, we initialize the early layers of the diffusion model with high-quality semantic features from pretrained models; in Phase 2, we transition to standard diffusion training with a gradually diminishing representation alignment term, allowing the model to increasingly focus on generation. This mirrors the sampling loop: first infer representation, then generate.
Phase 1: Representation Warmup Stage
@@
To alleviate the burden of learning semantic features from scratch, we begin with a dedicated warmup stage. During this phase, the model’s L2R circuit is initialized to align with semantically rich features extracted from a pretrained representation model (e.g., DINOv2, MAE, or CLIP). Let $\mathcal{H}_{\mathbf{\bm{\theta}}}(\mathbf{x})$ denote an encoder that maps an image $\mathbf{x}\in\mathcal{X}$ to its latent representation $\mathbf{z}\in\mathcal{Z}$, and let $\mathbf{f}_{\mathrm{rep}}:\mathcal{X}\to\mathcal{R}$ be a high-quality pretrained representation model. We use a single alignment objective shared by both phases:
$\mathcal{L}_{\mathrm{align}}(k)\,=\,\mathbb{E}_{\mathbf{x},\,\bm{\epsilon},\,t}\Big[\,s(k,t)\,\ell_{\mathrm{NT\text{-}Xent}}\Big(\mathcal{T}_{\bm{\theta}}(\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)),\;\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\Big)\,\Big]\,.$ (13)
Here $\mathbf{z}_{t}$ and the schedule $s(k,t)$ are
$\mathbf{z}_{t}\,=\,\begin{cases}\mathcal{H}_{\bm{\theta}}(\mathbf{x})&(t{=}0)\\ \alpha_{t}\,\mathbf{z}_{0}+\sigma_{t}\,\bm{\epsilon}&(t{>}0),\;\mathbf{z}_{0}=\mathcal{H}_{\bm{\theta}}(\mathbf{x})\end{cases}\quad\text{and}\quad s(k,t)\,=\,\begin{cases}1&(t{=}0)\\ \lambda_{\mathrm{train}}(k)=c_{0}\,\exp\!\left(-\frac{k}{\tau}\right)&(t{>}0)\,.\end{cases}$
Warmup sets t=0t{=}0; Phase 2 samples $t{\sim}\mathcal{U}[0,1]$ and uses the decayed $\lambda_{\mathrm{train}}(k){=}\,c_{0}\,\exp\!\big(-k/\tau\big)$ to gradually shift focus from alignment to generation, where $k$ is the training step, and $c_{0},\tau$ are hyperparameters.
Phase 2: Generative Training with Decaying Representation Guidance
@@
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{diffusion}}+\lambda_{\mathrm{train}}(k)\cdot\mathcal{L}_{\mathrm{align}}$ (14)
Here, $\mathcal{L}_{\mathrm{diffusion}}$ denotes the velocity prediction loss as defined in Eq. (5), and the alignment term is the objective in Eq. (13). The weight $\lambda_{\mathrm{train}}(k)$ modulates the impact of alignment during training. In practice, we instantiate $\ell_{\mathrm{NT\text{-}Xent}}$ with in-batch negatives and use the same projection head $\mathcal{T}_{\bm{\theta}}$ across both phases ( Section 4.1). The alignment thus acts as a weak semantic tether late in training, mitigating forgetting while letting R2G dominate. Both phases share the same alignment loss $\ell_{\mathrm{NT\text{-}Xent}}\big(\mathcal{T}_{\bm{\theta}}(\cdot),\mathbf{f}_{\mathrm{rep}}(\cdot)\big)$; they differ only in (i) the noise level of the input (clean t=0t{=}0 in Phase 1 vs. noisy t>0t{>}0 in Phase 2) and (ii) the schedule $\lambda_{\mathrm{train}}(k)$ (absent in warmup, exponentially decayed in Phase 2). Consistent with the augmented-space identity (Thm. 1), the surrogate gradient decomposes as
$\nabla_{\bm{\theta}}\,\mathcal{L}_{\mathrm{total}}(k)\;\approx\;\mathbb{E}\Big[\underbrace{\nabla_{\bm{\theta}}\,\mathcal{L}_{\mathrm{diffusion}}}_{{\color[rgb]{0.21875,0.3125,0.62890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.21875,0.3125,0.62890625}\textbf{\text{shapes}}}\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)}\;+\;s(k,t)\,\underbrace{\nabla_{\bm{\theta}}\,\mathcal{L}_{\mathrm{align}}(k)}_{{\color[rgb]{0.21875,0.3125,0.62890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.21875,0.3125,0.62890625}\textbf{\text{shapes}}}\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)}\Big]\,$,
up to standard surrogate mismatches. This makes the training-time decomposition mirror the sampling-time loop: first representation (L2R), then generation (R2G).